OpenStudy (anonymous):

I'm stuck on this problem, help! -0.5(x+8.5)=-0.3x-10.3 Someone please show me the steps

OpenStudy (anonymous):

-0.5*x-0.5*8.5=-0.3-10.3 -0.5x-4.25=-10.6 -0.5x=-10.6+4.25 -0.5x=-6.35 minus sign gets cancelled so, 0.5x=6.35 x=6.35/0.5 x=1.27

OpenStudy (anonymous):

thanks! but i wrote it wrong. its suppose to be Medals 0 -0.5(x+8.5)=-0.3x-10.3

OpenStudy (anonymous):

haha...okay -0.5*x-0.5*8.5=-0.3x-10.3 -0.5x-4.25=-0.3x-10.3 -0.5x+0.3x=-10.3+4.25 -0.2x=-6.05 minus sign gets cancelled so, 0.2x=6.05 x=6.05/0.2 x=30.25

OpenStudy (anonymous):

oh my goodness thank you!!!!!

OpenStudy (anonymous):

u welcm